OpenStudy (anonymous):

What is the 7th term of the geometric sequence 3, 9, 27, ...? 2,187 6,561 19,683 59,049

OpenStudy (lgbasallote):

\[\large A_n = A_1 r^{n-1}\] \[\large A_7 = 3 (3)^{7-1}\]

OpenStudy (anonymous):

the first 3^1 the second is 3^2 the 7th is 3^7

OpenStudy (anonymous):

3^7=2187

OpenStudy (lgbasallote):

3^7 o.O where did my solution go wrong sir @eliassaab ? @_@ i dont see where my analysis went wrong :(

OpenStudy (lgbasallote):

oh *facepalm* lol it's just the same haha